OpenStudy (sleepyjess):

Find f(x) and g(x) so the function can be expressed as y = f(g(x)). y = 7/x^2 + 10

OpenStudy (anonymous):

there are many things you could do depending on how creative you are try g(x) = 7/x^2

OpenStudy (sleepyjess):

What do you mean? I don't understand this.

OpenStudy (sleepyjess):

@hartnn

hartnn (hartnn):

there are several pairs of function you can make, let me give you an example : if y = 4/ log 3x +22 you can make f = 4/x +22 , g = log 3x which makes fog = 4/ log 3x +22 as required or you can make f = x +22 , g =4/ log 3x which also makes fog = 4/ log 3x +22 as required see if this makes sense

OpenStudy (sleepyjess):

Not really, is there anything I can do that doesn't use log?

hartnn (hartnn):

which function do you want to use ?

hartnn (hartnn):

if y = 4/ x^8+22 you can make f = 4/x +22 , g =x^8 which makes fog = 4/ x^8 +22 as required or you can make f = x +22 , g =4/ x^8 which also makes fog = 4/ x^8 +22 as required

OpenStudy (sleepyjess):

ok so would i just put 7/x^2 + 10 = 4/x^8 +22?

hartnn (hartnn):

omgee, 4/x^8 +22 is just an example :P

hartnn (hartnn):

what i did with 4/x^8 +22, try to do same thing with 7/x^2 + 10

hartnn (hartnn):

i just changed the numbers from your original question and solved it, so that you can "see" what i did and solve your question accordingly

OpenStudy (sleepyjess):

f=7/x^2 g=x+10 would that work?

hartnn (hartnn):

ofcourse! good :)

OpenStudy (sleepyjess):

so for the answer to explain it would i just put that adding them together makes 7/x^2 + 10

hartnn (hartnn):

f(x)=7/x^2 g(x)=x+10 f(g(x)) = (7/x^2) +10 as required.

hartnn (hartnn):

explain it like that ^^

OpenStudy (sleepyjess):

ok thank you so much!

hartnn (hartnn):

welcome ^_^